### v1

# anchor decoding AGC automatic gain control

[…]
AD
anchor decoding
AGC
automatic gain control
API
application program interface
AWGN
additive white Gaussian noise
AVX
advanced vector extensions
BCH
Bose–Chaudhuri–Hocquenghem
BDD
bounded distance decoding
BEC
binary erasure channel
BEE-PC
binary message passing based on error-and-erasure (EaE) decoding for product codes
BER
bit error rate
BDMC
binary DMC
BI-AWGN
binary input additive white Gaussian noise
BP
belief propagation
BPSK
binary phase shift keying
BSC
binary symmetric channel
CEL
Communications Engineering Lab
CPU
central processing unit
DFT
discrete Fourier transform
DMC
discrete nemoryless channel
DRS
dynamic reliability score
DRSD
dynamic reliability score (DRS) decoder
dSNR
design signal-to-noise ratio (SNR)
DSP
digital signal processing
DTP
decoding transition probability
EaE
error-and-erasure
EaED
EaE decoder
EEP
error-evaluator polynomial
ELP
error-locator polynomial
EMP
extrinsic message passing
MSP
minimal stall pattern
FEC
forward error correction
FER
frame error rate
FFT
fast Fourier transform
GPP
general purpose processor
GPC
generalized PC
GMD
generalized minimal distance
GMDD
generalized minimal distance decoder
HD
hard decision
HDD
hard decision decoding
HIHO
hard input hard output
HRB
highly reliable bit
iBDD
iterative bounded distance decoding (BDD)
IDFT
inverse discrete Fourier transform
iEaED
iterative error-and-erasure decoding
KIT
Karlsruhe Institute of Technology
LDPC
low-density parity-check
LR
likelihood ratio
LLR
log-likelihood ratio
LTE
Long Term Evolution
MAP
maximum a posteriori
MC
miscorrection
MD
miscorrection detection
MDS
maximum distance separable
ML
maximum likelihood
NCG
net coding gain
OSD
ordered statistics decoding
PC
product code
PCM
parity check matrix
PDF
probability density function
RS
Reed–Solomon
RV
random variable
SABM
soft-aided bit marking
SA-HDD
soft-aided hard decision decoding (HDD)
SABM-SR
SABM with scaled reliabilities
SDD
soft decision decoding
SISO
soft input soft output
SNR
signal-to-noise ratio
SPC
single parity check
TPD
turbo product decoding
On the Error Rate of Binary BCH Codes under Error-and-erasure Decoding
Sisi Miao, Jonathan Mandelbaum, Holger Jäkel, and Laurent Schmalen
Abstract
Determining the exact decoding error probability of linear block codes is an interesting problem. For binary BCH codes, McEliece derived methods to estimate the error probability of a simple bounded distance decoding (BDD) for BCH codes. However, BDD falls short in many applications. In this work, we consider error-and-erasure decoding and its variants that improve upon BDD. We derive closed-form expressions for their error probabilities and validate them through simulations. Then, we illustrate their use in assessing concatenated coding schemes.
I Introduction
Binary Bose–Chaudhuri–Hocquenghem (BCH) codes with BDD are widely used in communication systems [1]. To meet the growing performance demands, enhancing their coding gain through soft-decision decoding has become important. Decoders such as Chase II [2] and ordered statistics decoding (OSD[3] can approach maximum likelihood (ML) performance but at prohibitive complexity. In this paper, we consider EaE decoding, perhaps the simplest soft-decision decoder for BCH codes. Apart from the hard-decision values, EaE decoding incorporates soft information by adding a single reliability class of erasures ($\mathord{?}$”) [4]. There exist different variants of EaE decoding. In particular, we focus on the decoder (EaED) constructed from two BDD steps [5, Sec. 3.8.1]. This algorithm offers good decoding performance and is closely related to Chase II decoding. Moreover, with a simple modification, it can reproduce the decoding result of Forney’s algebraic EaE decoder [6, 7].
One appealing property of BDD is that its decoding error probability can be determined analytically [8]: given the number of input errors, the distribution of residual errors after decoding is known. In contrast, no such results exist yet for EaE decoding. In this paper, we derive a closed-form expression for the considered EaED. Specifically, we aim to determine the decoding transition probability (DTP) P(R=r|U=u,E=e){P(R\!=\!r|U\!=\!u,E\!=\!e)}, where $R$ is the number of residual errors after decoding, $U$ the number of input errors, and $E$ the number of input erasures. With the DTP, the post-decoding bit error rate (BER) and frame error rate (FER) for various memoryless channels can be computed. The DTP is validated through simulations and applied in a concatenated coding scheme.
[…]
II Preliminaries
Let $\mathcal{C}[n,k,d_{\textup{des}}]$ be a primitive, narrow-sense, binary BCH code, where $n$ is the block length, $k$ the dimension, and $d_{\textup{des}}$ the designed distance, a lower bound on the minimum distance $d_{\textnormal{min}}$. The code rate is given by $R_{c}=k/n$. Let $t$ denote the number of correctable errors. The parameters are related by $n=2^{b}-1$, $k\geq n-bt$, and $2t+1=d_{\textup{des}}$. Let $A_{w}$ denote the number of codewords of weight $w$ in $\mathcal{C}$. Binary primitive BCH codes have approximately binomial weight distributions given by $A_{w}\approx 2^{-bt}\binom{n}{w}$ for $d_{\mathrm{des}}\leq w\leq n-d_{\mathrm{des}}$, with $A_{0}=A_{n}=1$, and $A_{w}=0$ elsewhere [1]. The exact weight enumerators of some short BCH codes are listed in [9].
We consider three communication channels: the binary symmetric channel (BSC) with cross-over probability $\delta=P(Y\neq X)$}, the EaE channel, and the binary input additive white Gaussian noise (BI-AWGN) channel. In the EaE channel, the input is $X\in\mathbb{F}_{2}$} while the output is $Y\in\{0,1,\mathord{?}\}$}. The channel transition probabilities are $P(Y=1-x|X=x)=\delta_{c}$}, $P(Y=\mathord{?}|X=x)=\epsilon_{c}$}, and $P(Y=x|X=x)=1-\delta_{c}-\epsilon_{c}$}. For the BI-AWGN channel, we map bit $x_{i}\mapsto(-1)^{x_{i}}=:\tilde{x}_{i}$}, and the received symbol is $\tilde{y}_{i}=\tilde{x}_{i}+n_{i}\in\mathbb{R}$}, where $n_{i}\sim\mathcal{N}(0,\sigma_{n}^{2})$} with $\sigma_{n}^{2}=(2R_{c}E_{b}/N_{0})^{-1}$. In the following, by introducing a threshold $\mathsf{T}\in\mathbb{R}$}, we reduce the BI-AWGN to an EaE channel: if $\tilde{y}_{i}\in[-\mathsf{T},\mathsf{T}]$}, an erasure is declared, i.e., $y_{i}=\mathord{?}$, otherwise $y_{i}=\mathrm{sign}(\tilde{y}_{i})$}. Thus, the parameters of the resulting EaE are $\delta_{c}=Q\left(\tfrac{\mathsf{T}+1}{\sigma}\right)$, and $\epsilon_{c}=1-Q\left(\tfrac{\mathsf{T}-1}{\sigma}\right)-\delta_{c}.$ For $\mathsf{T}=0$, we recover a BSC with $\delta=Q\left(\tfrac{1}{\sigma}\right)$.
We use $\bm{x}\in\mathcal{C}$ to denote the transmitted codeword and $\bm{y}$ to denote the received word, whose alphabet is determined by the channel. At the output of a BSC, let $\mathcal{S}_{t}(\bm{c})=\{\bm{y}\in\mathbb{F}_{2}^{n}:d(\bm{y},\bm{c})\leq t\}$ denote the Hamming sphere of radius $t$ for $\bm{c}\in\mathcal{C}$. Then
$\textsf{BDD}\left(\bm{y}\right)=\begin{cases}\bm{c},&\text{if }\bm{y}\in\mathcal{S}_{t}(\bm{c}),\\ \bm{y}\,(\mathsf{fail}),&\text{otherwise}.\end{cases}$
Note that the decoding outcome falls into one of three categories: (i) decoding success, if $\bm{c}=\bm{x}$; (ii) decoding failure if no codeword is found; or (iii) miscorrection (also called a decoding error as in [8]), if $\bm{c}\neq\bm{x}$.
For each decoder considered in this paper, we distinguish these three cases when computing the DTP. Taking BDD as an example, in addition to the random variables $U$ and $R$, we introduce the RV $D\in\mathcal{D}:=\{\mathsf{succ},\mathsf{fail},\mathsf{mc}\}$} representing the decoding outcome for success, failure, and miscorrection, respectively. Then, we define the conditional joint probability $P_{\textnormal{{BDD}}}(R=r,D=d|U=u)$}, i.e., the probability of observing outcome $d$ with $r$ residual errors given $u$ input errors. For simplicity, we use the shorthand notation $P_{\textnormal{{BDD}}}^{d}(r|u):={P_{\textnormal{{BDD}}}(r,d|u)}$, yielding $P_{\textnormal{{BDD}}}^{\textnormal{succ}}(r|u)$, $P_{\textnormal{{BDD}}}^{\textnormal{fail}}(r|u)$, and $P_{\textnormal{{BDD}}}^{\textnormal{mc}}(r|u)$. It follows that
$P_{\textnormal{{BDD}}}(r|u)=P_{\textnormal{{BDD}}}^{\textnormal{mc}}(r|u)+P_{\textnormal{{BDD}}}^{\textnormal{fail}}(r|u)+P_{\textnormal{{BDD}}}^{\textnormal{succ}}(r|u).$
Note that for a decoding failure, $P_{\textnormal{{BDD}}}^{\textnormal{fail}}(R\neq u|u)=0$, since $\textsf{BDD}\left(\bm{y}\right)=\bm{y}$ in this case, and for a decoding success, $P_{\textnormal{{BDD}}}^{\textnormal{succ}}(R\neq 0|u)=0$. Furthermore, we denote by $P_{\textnormal{{BDD}}}^{\textnormal{succ}}(u)$, $P_{\textnormal{{BDD}}}^{\textnormal{fail}}(u)$, and $P_{\textnormal{{BDD}}}^{\textnormal{mc}}(u)$ the conditional probabilities of success, failure, and miscorrection given $u$ input errors, respectively.
Since we study linear codes with a symmetric decoding algorithm, we assume the transmission of $\bm{x}=\bm{0}$} throughout the paper. Then, the decoder input errors are $u=\textnormal{wt}\left(\bm{y}\right)$}.
III Analysis of Bounded Distance Decoding
[…]
Figure 1: Graphical illustration of the miscorrection scenario when using a BDD.
In this section, we consider BDD after BSC where the received sequence is $\bm{y}\in\mathbb{F}_{2}^{n}$, i.e., without erasures. To this end, we review the method in [8] and refine it slightly such that it also gives the detailed DTP for BDD.
Case 1: If $U=u\leq t$, then $P_{\textnormal{{BDD}}}^{\textnormal{succ}}(0|u)=1$ and $P_{\textnormal{{BDD}}}^{d}(r|u)=0$ for $d\neq\mathsf{succ},r\neq 0$.
Case 2: If U=u>t{U=u>t}, either a miscorrection or a decoding failure may occur. In the miscorrection scenario, we have $P_{\textnormal{{BDD}}}^{\textnormal{mc}}(r|u)=N(u)/\tbinom{n}{u}$, where $N(u)$ is the number of received patterns $\bm{y}$ with $\textnormal{wt}\left(\bm{y}\right)=u$ that satisfy $\bm{y}\in\mathcal{S}_{t}(\bm{c})$ for some $\bm{c}\in\mathcal{C}$ with $\textnormal{wt}\left(\bm{c}\right)=r$. Figure 1 illustrates this scenario with the support of $\bm{c}$ permuted to the right.
We define $a=|\{i\in[n]:c_{i}=1,y_{i}=0\}|$} as the number of correct bits in $\bm{y}$ that are (wrongly) flipped by BDD, and $b=|\{i\in[n]:c_{i}=0,y_{i}=1\}|$} as the number of erroneous positions in $\bm{y}$ that are (correctly) flipped by BDD. Hence, it holds that u+ab=r{u+a-b=r} and $a+b=d\left(\bm{y},\bm{c}\right)\leq t$. For fixed $a$ and $b$, there are $\binom{r}{a}\binom{n-r}{b}=\binom{r}{a}\binom{n-r}{u+a-r}$ distinct patterns for $\bm{y}$. Additionally, the number of codewords $\bm{c}\in\mathcal{C}$ with $\textnormal{wt}\left(\bm{c}\right)=r,$ is given by the weight enumerator $A_{r}$. Thus, we obtain
$P_{\textnormal{{BDD}}}^{\textnormal{mc}}(r|u)=\frac{N(u)}{\tbinom{n}{u}}=\frac{\sum_{(a,b)\in\mathcal{S}_{a,b}}A_{r}\tbinom{r}{a}\tbinom{n-r}{b}}{\tbinom{n}{u}}$.
where $\mathcal{S}_{a,b}=\{(a,b)\in[t]\times[t]:a+b\leq t,u+a-b=r\}$}. Then, as $P_{\textnormal{{BDD}}}^{\textnormal{mc}}(r|u)\!+\!P_{\textnormal{{BDD}}}^{\textnormal{fail}}(r|u)\!=\!1$}, we have
$P_{\textnormal{{BDD}}}^{\textnormal{fail}}(r|u)=1-\sum_{\tilde{r}=u-t}^{u+t}P_{\textnormal{{BDD}}}^{\textnormal{mc}}(\tilde{r}|u).$
The results are summarized in Theorem 1.
Theorem 1.
The DTP of BDD is given as follows. For $u\leq t$, $P_{\textnormal{{BDD}}}(0|u)=1$, $P_{\textnormal{{BDD}}}(r|u)=0$, $(d,r)\neq(\mathsf{succ},0)$. For $u>t$,
$P_{\textnormal{{BDD}}}(r|u)\!\!=\!\!\begin{cases}P_{\textnormal{{BDD}}}^{\textnormal{mc}}(r|u),&\!\!\!u-t\!\leq\!r\!\leq\!u+t,\!r\!\neq\!u,\\ P_{\textnormal{{BDD}}}^{\textnormal{mc}}(r|u)\!+\!P_{\textnormal{{BDD}}}^{\textnormal{fail}}(r|u),&\!\!\!r=u,\\ 0,&\!\!\!\text{otherwise.}\end{cases}$
IV Error-and-erasure Decoding
IV-A Error-and-erasure (EaE) Decoding Algorithm
In the following, we revisit the EaED algorithm presented in [10, Algorithm 2]. For a received word $\bm{y}\in\{0,\mathord{?},1\}^{n}$ with $E$ erasures and $U$ errors, let $\bm{w}:=\textnormal{EaED}(\bm{y})$}.
In EaED, a pair of random filling patterns $(\bm{p}^{(1)},\bm{p}^{(2)})$ with $\bm{p}^{(1)},\bm{p}^{(2)}\in\mathbb{F}_{2}^{E}$} and $\bm{p}^{(1)}+\bm{p}^{(2)}=\bm{1}_{E}$} is constructed, by drawing $\bm{p}^{(1)}$ as a length-$E$ Bernoulli($\frac{1}{2}$) vector. The erasures in $\bm{y}$ are then replaced by $\bm{p}^{(1)}$ and $\bm{p}^{(2)}$ to form the test patterns $\bm{y}^{(1)},\bm{y}^{(2)}\in\mathbb{F}_{2}^{n}$. Two BDDs yield $\bm{w}^{(i)}:=\textsf{BDD}\left(\bm{y}^{(i)}\right)$, $i\in\{1,2\}$. The final output $\bm{w}$ is selected from $\bm{w}^{(1)}$ and $\bm{w}^{(2)}$ according to the following rules:
Case 1: If both BDD fail, set $\bm{w}=\bm{y}$ and declare a failure.
Case 2: If $\bm{w}^{(i)}\in\mathcal{C}$ for exactly one $\bm{w}^{(i)}$, set $\bm{w}=\bm{w}^{(i)}$.
Case 3: If both BDD succeed, let $d_{i}=d_{\sim{\textnormal{E}(\bm{y})}}\left(\bm{w}^{(i)}\right)$} for $i\in\{1,2\}$}. Then, choose $\bm{w}=\bm{w}^{(1)}$} if $d_{1}<d_{2}$}, and $\bm{w}=\bm{w}^{(2)}$} if $d_{2}<d_{1}$. If $d_{1}=d_{2}$, the output $\bm{w}$ is chosen as $\bm{w}^{(1)}$ or $\bm{w}^{(2)}$ with equal probability.
Remark 1: To avoid miscorrections, one commonly declares a decoding failure if the number of erasures $E$ exceeds a certain value before decoding. This can be easily incorporated in the upcoming DTP derivation by truncation.
Remark 2: The EaED reverts to a conventional BDD when there are no erasures, i.e., $e=0$.
Remark 3: For simplicity, when EaED declares a decoding failure, each residual erasure in $\bm{y}$ is resolved by an independent Bernoulli($\frac{1}{2}$) trial, such that on average every residual erasure contributes to $\frac{1}{2}$ bit error.
IV-B Decoding Transition Probability
Theorem 2 outlines our derived $P_{\textnormal{{EaED}}}(r|u,e)$ of EaED.
Theorem 2.
The DTP of EaED is given as follows. For $2u+e<d_{\textnormal{min}}$}, $P_{\textnormal{{EaED}}}(0|u,e)=1$}. For $2u+e\geq d_{\textnormal{min}}$}, $P_{\textnormal{{EaED}}}(r|u,e)=\sum_{d\in\mathcal{D}}P_{\textnormal{{EaED}}}^{d}(r|u,e)$ and
$P_{\textnormal{{EaED}}}^{d}(r|u,e)=\textstyle\sum_{e_{1}=0}^{e}P_{\textnormal{{EaED}}}^{d}(r|u,e,e_{1})\cdot P(e_{1}|u,e),$
where $e_{1}=\!\textnormal{wt}\left(\bm{p}^{(1)}\right)$}. The value $P_{\textnormal{{EaED}}}^{d}(r|u,e,e_{1})$ is given by (4) when $e_{1}\in[0,t-u]$}. For $e_{1}\in[u+e-t,e]$}, we have the symmetry relation $P_{\textnormal{{EaED}}}^{d}(r|u,e,e_{1})=P_{\textnormal{{EaED}}}^{d}(r|u,e,e\!-\!e_{1})$}. For $e_{1}\in(t-u,u+e-t)$}, it is approximated by (5) and (A).
Proof.
See Appendix A. ∎
Note that, due to the Bernoulli trials to resolve residual erasures upon decoding failure (Remark 3), $R$ can now take values in $\{0,\tfrac{1}{2},1,\tfrac{3}{2},\dots,n\}$.
TABLE I: Probability of success of EaED with $u$ errors and $e$ erasures for the $[255,239,5]$ primitive BCH code.
$u$ $e$ 0 1 2 3 4 5 6 7 8
0 1.000 1.000 1.000 1.000 1.000 0.999992 0.688 0.453 0.289
1 1.000 1.000 1.000 0.998 0.622 0.371 0.216 0.123 0.069
2 1.000 0.753 0.376 0.186 0.093 0.046 0.023 0.012 0.006
TABLE II: Probability of miscorrection of EaED with $u$ errors and $e$ erasures for the $[255,239,5]$ primitive BCH code.
$u$ $e$ 0 1 2 3 4 5 6 7 8
0 0 0 0 0 0 7.8e-06 0.233 0.407 0.530
1 0 0 0 0.002 0.282 0.469 0.585 0.655 0.695
2 0 0.247 0.497 0.622 0.684 0.716 0.732 0.740 0.744
3 0.494 0.744 0.745 0.746 0.747 0.747 0.748 0.748 0.748
4 0.494 0.746 0.747 0.748 0.748 0.748 0.748 0.748 0.748
5 0.498 0.748 0.748 0.748 0.748 0.748 0.748 0.748 0.748
[…]
Figure 2: Simulated and computed BER results of the $[255,239,5]$ primitive BCH code. The $\mathsf{EaED}$ uses the parameter $\mathsf{T}=0.16$ and the ${\textnormal{{EaED}}}^{\textsf{a}}$ uses the parameters $\mathsf{T}=0.13$ and $\mathsf{T}_{\textnormal{a}}=0.75$. The OSD uses order 2, as higher orders do not yield additional performance gains.
Using the $[255,239,5]$ BCH code as an example, Tab. I and Tab. II list the conditional probabilities $P_{{\textnormal{{EaED}}}}^{\textnormal{succ}}(u,e)$ and $P_{{\textnormal{{EaED}}}}^{\textnormal{mc}}(u,e)$, respectively, which are calculated using (2) and marginalization over $U$. Note that if $e=0$, Theorem 2 reverts to Theorem 1. Accordingly, the first column of both tables corresponds to BDD. The miscorrection rate of BDD is high and is close to the upper bound $\frac{1}{t!}$ derived in [8]. Compared with BDD, EaED can decode beyond the minimum distance. When $2u+e$ slightly exceeds $d_{\textnormal{min}}-1$, the decoding success rate of EaED is still high. However, this comes at the expense of an increased probability of miscorrection for larger values of $2u+e$. The post-decoding BERs of EaED and BDD are compared in Fig. 2 with the threshold $\mathsf{T}$ optimized to minimize the BER of EaED at $9$ dB. We see that the simulation and computed values closely agree. Additionally, as expected, EaED improves upon BDD in terms of performance.
V Miscorrection Reduction with Anchor Bits
Miscorrections cause undetectable errors that can be more harmful than decoding failures, e.g., when retransmission is allowed upon decoding failure, or during iterative decoding of product codes with BCH component codes, where miscorrections lead to error propagation.
To mitigate miscorrections, we consider anchor bits [11, 10], which form an additional reliability class. In contrast to erasures, anchor bits are assumed to be highly reliable, and any decoding decision that conflicts with them is treated as a miscorrection. Let $\mathcal{H}_{\mathsf{a}}\subseteq[n]$ denote the indices of anchor bits. For a decoding result $\bm{w}=\mathsf{Dec}(\bm{y})$ with $\bm{w}\in\mathcal{C}$, we perform a miscorrection detection (MD) step: if there exists $i\in\mathcal{H}_{\mathsf{a}}$ such that $w_{i}\neq y_{i}$, the estimate $\bm{w}$ is rejected and the decoder outputs a failure. It remains to define the set of anchor bits $\mathcal{H}_{\mathsf{a}}$. In this work, anchor bits are obtained by applying a threshold $\mathsf{T}_{\textnormal{a}}$ to the BI-AWGN channel output, i.e., by marking bits with output value $|\tilde{y}_{i}|>\mathsf{T}_{\textnormal{a}}$. Therefore, $\mathcal{H}_{\mathsf{a}}=\{i\in[n]:|\tilde{y}_{i}|>\mathsf{T}_{\textnormal{a}}\}$. Next, we show that the framework summarized in Theorems 1 and 2 can be extended to this scenario with minor modifications.
V-A BDD with Anchor Bits
In this section, we present the DTP of BDD aided by anchor bits, termed a ${\textnormal{{BDD}}}^{\textsf{a}}$ decoder. We define
$P_{\textnormal{ca}}=P(\tilde{Y}>\mathsf{T}_{\textnormal{a}}|\tilde{Y}>0,X=1)=Q\!\left(\tfrac{\mathsf{T}_{\textnormal{a}}-1}{\sigma}\right)/(1-\delta),$
$P_{\textnormal{wa}}=P(\tilde{Y}<-\mathsf{T}_{\textnormal{a}}|\tilde{Y}<0,X=1)=Q\!\left(\tfrac{-\mathsf{T}_{\textnormal{a}}-1}{\sigma}\right)/\delta$
to be the probability of a correct and erroneous anchor bit, respectively. Following the analysis of Sec. 1 and the illustration in Fig. 1, a miscorrection $\textsf{BDD}\left(\bm{y}\right)=\bm{c}$ is rejected by the anchor bits with probability $(1-P_{\textnormal{ca}})^{a}(1-P_{\textnormal{wa}})^{b}$. Thus,
$P_{{\textnormal{{BDD}}}^{\textsf{a}}}^{\textnormal{mc}}(r|u)=\sum_{(a,b)\in\mathcal{S}_{a,b}}A_{r}\tbinom{r}{a}\tbinom{n-r}{b}(1-P_{\textnormal{ca}})^{a}(1-P_{\textnormal{wa}})^{b}/\tbinom{n}{u}$,
where $(1-P_{\textnormal{ca}})^{a}\ll(1-P_{\textnormal{wa}})^{b}$ since typically $P_{\textnormal{ca}}\gg P_{\textnormal{wa}}$. $P_{{\textnormal{{BDD}}}^{\textsf{a}}}^{\textnormal{fail}}$ and $P_{{\textnormal{{BDD}}}^{\textsf{a}}}^{\textnormal{succ}}$ can then be calculated easily.
V-B EaED with Anchor Bits
Similar to BDDa, we introduce EaEDa, which is an EaED with the MD step using anchor bits, whose DTP is denoted by $P_{{\textnormal{{EaED}}}^{\textsf{a}}}(r|u,e)$. To be precise, in EaEDa, for $\bm{w}^{(i)}=\textsf{BDD}\left(\bm{y}^{(i)}\right),i\in\{1,2\}$, we check if there exists $j\in\mathcal{H}_{\mathsf{a}}$ such that $w^{(i)}_{j}\neq y_{j}$. If this is true, we declare a decoding failure for $\textsf{BDD}\left(\bm{y}^{(i)}\right)$. Then, we proceed to select the final output as described in Sec. IV-A.
Note that we always set $\mathsf{T}_{\textnormal{a}}>\mathsf{T}$ to ensure that erasures cannot be anchor bits. In this case, $P_{\textnormal{ca}}$ and $P_{\textnormal{wa}}$ are given as
$P_{\textnormal{ca}}=P(\tilde{Y}>\mathsf{T}_{\textnormal{a}}|\tilde{Y}>\mathsf{T},X=1)=Q\left(\tfrac{\mathsf{T}_{\textnormal{a}}-1}{\sigma}\right)/(1-\delta_{c}-\epsilon_{c}),$
$P_{\textnormal{wa}}=P(\tilde{Y}<\mathsf{T}_{\textnormal{a}}|\tilde{Y}<\mathsf{T},X=1)=Q\left(\tfrac{-\mathsf{T}_{\textnormal{a}}-1}{\sigma}\right)/\delta_{c}.$
The DTP of EaEDa is calculated in Appendix B.
V-C Improving one-shot decoding of BCH code
As shown in Fig. 2, EaEDa outperforms both EaED and BDD. Next, we examine the miscorrection rate. For EaED, most bit errors are due to miscorrections, whereas BDD exhibits a smaller fraction. This observation aligns with Tab. I and Tab. II, reflecting that EaED has a stronger error correction capability but is more prone to miscorrection. By contrast, EaEDa substantially reduces miscorrections through the use of anchor bits. Thus, EaEDa improves decoding performance while maintaining a low miscorrection rate, keeping the undetectable error rate negligible.
[…]
Figure 3: FER versus crossover probability for eight RS[544,514,15][544,514,15] outer code interleaved with 6464 BCH[700,680,2][700,680,2] inner code. The EaEDa uses the parameters $\mathsf{T}=0.05$ and $\mathsf{T}_{\textnormal{a}}=0.56$ which minimizes the FER at $7$ dB.
VI Application in Concatenated BCH-RS Schemes
In [12], a concatenated scheme with Reed–Solomon (RS) outer codes and BCH inner codes is presented and thoroughly analyzed. In this section, we show that only replacing the BDD used for the BCH codes in that scheme with EaEDa lowers the post- forward error correction (FEC) FER with only a small increase in complexity. The rest of the scheme follows [12].
The scheme in [12] uses shortened BCH codes to reduce miscorrections and match the target code rate. Let $\mathcal{H}_{\mathsf{s}}=[n_{\text{short}]}$} denote the set of shortened positions. Then, the subcode
$\mathcal{C}^{n}_{\text{short}}:=\{\bm{c}\in\mathcal{C}[n,k]:c_{i}=0,\forall i\in\mathcal{H}_{\mathsf{s}}\}\subseteq\mathcal{C}[n,k]$
is isomorphic to the shortened code $\mathcal{C}_{\text{short}}[\,n-n_{\text{short}},\,k-n_{\text{short}}]$ obtained by neglecting positions in $\mathcal{H}_{\mathsf{s}}$. Note that shortened bits are not transmitted, such that $\mathcal{C}_{\text{short}}$ is used for transmission and $\mathcal{H}_{\mathsf{a}}\cap\mathcal{H}_{s}=\emptyset$. We further assume that the error locations that the BDD declares are randomly distributed, and therefore, the rate that such an error location is in $\mathcal{H}_{\mathsf{s}}$ is $n_{\text{short}}/n$, which provides a good approximation when $n_{\text{short}}$ is small [12]. Therefore, we treat the shortened bits as additional correct anchor bits after decoding on $\mathcal{C}$. We can then compute the DTP of EaEDa using the method derived in Appendix B.
[…]
VII Conclusion
In this paper, we derive a closed-form expression for estimating the decoding behavior of EaED. Furthermore, results are extended to incorporate an enhanced version of the decoder. The MATLAB scripts used to generate the results are available at https://github.com/kit-cel/EaED_BCH.
Appendix A EaED Calculation
TABLE III: All possible BDD outcomes when decoding the two test patterns in EaED, assuming that $2e+u\geq d_{\textnormal{min}}$.
$\textsf{BDD}\left(\bm{y}^{(1)}\right)$ $\bm{0}$ $\bm{0}$ $\mathsf{fail}$ $\bm{c}$ $\mathsf{fail}$ $\bm{c}$ $\bm{c}$ $\bm{c}$ $\mathsf{fail}$
$\textsf{BDD}\left(\bm{y}^{(2)}\right)$ $\bm{c}$ $\mathsf{fail}$ $\mathsf{fail}$ $\mathsf{fail}$ $\bm{c}$ $\bm{c}$ $\bm{c}^{\prime}$ $\bm{0}$ $\bm{0}$
[…]
Figure 4: Sorting of the cases in Tab. III according to the value of $e_{1}$. When $u>t$, sets $\mathcal{L}$ and $\mathcal{R}$ will not occur.
[…]
Figure 5: Graphical illustration of case ➁. Orange parts sum up to $e_{1}$ while green parts sum up to $e_{2}=e-e_{1}$. The gray parts sum up to $d_{\sim{\textnormal{E}(\bm{y})}}(\bm{c})$.
In this appendix, we derive DTP $P_{\mathsf{EaED}}(r|u,e)$ of EaED.
Case 1: If $e=0$, EaED reverts to BDD. Therefore, the transition probabilities are given by Theorem 1.
Case 2: Consider $e\neq 0$.
Case 2.1: If $2u+e<d_{\textnormal{min}}$, $P_{{\textnormal{{EaED}}}}^{\textnormal{succ}}(u,e)=1$ [7].
Case 2.2: When $2u+e\geq d_{\textnormal{min}}$}, EaED may still succeed. Hence, we must calculate $P_{{\textnormal{{EaED}}}}^{\textnormal{mc}}(r|u,e)$}, $P_{{\textnormal{{EaED}}}}^{\textnormal{fail}}(r|u,e)$}, and $P_{{\textnormal{{EaED}}}}^{\textnormal{succ}}(r|u,e)$} separately for this case.
Let the RV $E_{1}=\textnormal{wt}\left(\bm{p}^{(1)}\right)$; under the all-zero codeword assumption, $E_{1}=e_{1}$ corresponds to erroneous bits introduced by the random filling pattern to $\bm{y}^{(1)}$ before BDD, i.e., $\textnormal{wt}\left(\bm{y}^{(1)}\right)=u+e_{1}$. Analogously, define $e_{2}=\textnormal{wt}\left(\bm{p}^{(2)}\right)$. Note that $E_{1}+E_{2}=E$}. Consequently, $\textnormal{wt}\left(\bm{y}^{(2)}\right)=u+e-e_{1}$}.
Table III lists all nine possible joint BDD outcomes. In particular, note that $\textsf{BDD}\left(\bm{y}^{(1)}\right)=\textsf{BDD}\left(\bm{y}^{(2)}\right)=\bm{0}$} cannot occur, since this would imply $u+e_{1}\leq t$ and $u+e-e_{1}\leq t$, which together yield $2u+e\leq 2t\leq d_{\textnormal{min}}-1$, contradicting $u+2e\geq d_{\textnormal{min}}$. Therefore, at most one of the two BDD can correctly decode to $\bm{0}$.
As depicted in Fig. 4, based on the value of $e_{1}$, the nine cases can be grouped into three categories (sets), denoted $\mathcal{L}:=\{$➀, ➁}\}, $\mathcal{M}=\{$➂, ➃, ➄, ➅, ➆}\}, and $\mathcal{R}=\{$➇, ➈}\}, corresponding to the cases $e_{1}\in[0,t-u]$, $e_{1}\in(t-u,u+e-t)$ and $e_{1}\in[u+e-t,e]$, respectively. Then, we can compute
$P_{{\textnormal{{EaED}}}}^{\textnormal{mc}}(r|u,e)=\textstyle\sum_{e_{1}=0}^{e}P_{{\textnormal{{EaED}}}}^{\textnormal{mc}}(r|u,e,e_{1})\cdot P(E_{1}=e_{1}|u,e)$
$=\underbrace{\!2\!\sum_{e_{1}=0}^{t-u}P_{{\textnormal{{EaED}}}}^{\textnormal{mc}}(r|u,e,e_{1})\frac{\binom{e}{e_{1}}}{2^{e}}}_{\mathrm{scenarios}\,\mathcal{L,R}}+\!\!\!\!\!\underbrace{\sum_{e_{1}=t-u+1}^{u+e-t-1}\!\!\!\!P_{{\textnormal{{EaED}}}}^{\textnormal{mc}}(r|u,e,e_{1})\frac{\binom{e}{e_{1}}}{2^{e}}}_{\mathrm{scenario}\,\mathcal{M}}.$ (1)
The factor $2$ before the first sum is due to the symmetry between $\mathcal{L}$ and $\mathcal{R}$. Next, we compute the probabilities associated with $\mathcal{L}$ and $\mathcal{M}$.
The set $\mathcal{L}$ contains the joint decoding outputs for which $e_{1}\leq t-u$, which implies that $\textsf{BDD}\left(\bm{y}^{(1)}\right)=\bm{0}$. Furthermore, $\textnormal{wt}\left(\bm{y}^{(2)}\right)=u+e-e_{1}\geq 2u+e-t\geq d_{\textnormal{min}}-t>t$. Therefore, $\textsf{BDD}\left(\bm{y}^{(2)}\right)$ is either a miscorrection (case ➀) or decoding failure (case ➁).
There are $\binom{n}{u}\binom{n-u}{e}\binom{e}{e_{1}}:=\Theta$} possible combinations of locations for the $u$ errors, $e$ erasures, and $e_{1}$ 11s among the $e$ erased positions when generating the test pattern $\bm{y}^{(1)}$. Then, we first count the number of events leading to ➀, from which ➁ follows directly.
Now, for case ➀, we fix a non-zero codeword $\bm{c}$ and first enumerate all possible $\bm{y}^{(2)}$. Based on each $\bm{y}^{(2)}$, we then infer and count all elementary events $\bm{y}$ consistent with the realization $(u,e,e_{1})$.
We now first count the number of $\bm{y}^{(2)}$ patterns with weight $u+e_{2}$ that fulfill $\bm{y}^{(2)}\in\mathcal{S}_{t}(\bm{c})$ for some $\bm{c}\in\mathcal{C}\setminus\{\bm{0}\}$}, with $\textnormal{wt}\left(\bm{c}\right)=r$. To this end, as illustrated in Fig. 5 and similarly to BDD, we first fix $\bm{y}^{(2)}$ and introduce variables $a=|\{i\in[n]:c_{i}=1,y^{(2)}_{i}=0\}|$} and $b=|\{i\in[n]:c_{i}=0,y^{(2)}_{i}=1\}|$}. As $\textsf{BDD}\left(\bm{y}^{(2)}\right)=\bm{c}$, $(a,b)\in\mathcal{S}_{a,b}=\{(a,b)\in[t]\times[t]:a+b\leq t,u+e\!-\!e_{1}+a-b\!=\!r\}$. There are $A_{r}\binom{r}{a}\binom{n-r}{b}$ possible such $\bm{y}^{(2)}$ patterns.
Then, to obtain the number of possible $\bm{y}$ patterns with $\bm{y}^{(2)}$ fixed, we distribute $e$ erasures across the four different parts of $\bm{y}^{(2)}$, which are divided according to the four possible values of $(c_{i},y^{(2)}_{i})$. Next, we introduce the variables $\lambda:=|\{i\in[n]:c_{i}=0,\,y_{i}=\mathord{?},\,y^{(1)}_{i}=1\}|$} and $\gamma:=|\{i\in[n]:c_{i}=0,\,y_{i}=\mathord{?},\,y^{(1)}_{i}=0\}$}. For a fixed pair of $\bm{c}$ and $\bm{y}^{(2)}$, there are $\binom{n-r-b}{\lambda}\binom{b}{\gamma}\binom{r-a}{e-e_{1}-\gamma}\binom{a}{e_{1}-\lambda}$ distinct sequences $\bm{y}$ when considering the possible locations of the erasures. Observe that whenever $\bm{y}^{(2)}$ and $\bm{y}$ are specified, the word $\bm{y}^{(1)}$ is uniquely determined. Hence, the number of placements for case ➀ amounts to
$\sum_{(a,b)\in\mathcal{S}_{a,b}}\sum_{\lambda\in\mathcal{S}_{\lambda}}\sum_{\gamma\in\mathcal{S}_{\gamma}}V(a,b,\lambda,\gamma)$ (2)
where $\mathcal{S}_{\lambda}=\{\lambda\in[n-r-b]:0\leq e_{1}-\lambda\leq a\}$, $\mathcal{S}_{\gamma}=\{\gamma\in[b]:0\leq e-e_{1}-\gamma\leq r-a\}$, and
$V(a,b,\lambda,\gamma)=A_{r}\tbinom{r}{a}\tbinom{n-r}{b}\tbinom{n-r-b}{\lambda}\tbinom{b}{\gamma}\tbinom{r-a}{e-e_{1}-\gamma}\tbinom{a}{e_{1}-\lambda}$,
which is abbreviated as $V$ when unambiguous.
Next, for case ➀, we show that evaluating (2) in a refined range of $\lambda$ and $\gamma$ yields the number of elementary cases that lead to different EaED outcomes.
As depicted in Fig. 5, $d_{\sim{\textnormal{E}(\bm{y})}}(\bm{c})=b-\gamma+a-(e_{1}-\lambda)$ and obvisouly $d_{\sim{\textnormal{E}(\bm{y})}}(\bm{0})=u$. Hence, if $d_{\sim{\textnormal{E}(\bm{y})}}(\bm{c})<d_{\sim{\textnormal{E}(\bm{y})}}(\bm{0})$}, a miscorrection occurs as EaED chooses $\bm{c}$ as its estimate. Similarly, when ${d_{\sim{\textnormal{E}(\bm{y})}}(\bm{c})}>d_{\sim{\textnormal{E}(\bm{y})}}(\bm{0})$, EaED chooses $\bm{0}$ as its output and will decode successfully. Finally, When $d_{\sim{\textnormal{E}(\bm{y})}}(\bm{c})=d_{\sim{\textnormal{E}(\bm{y})}}(\bm{0})$}, miscorrection and success happen with equal probability.
For convenience, let $\mathcal{S}_{\odot}:=\{(\lambda,\gamma):\lambda-\gamma\odot u+e_{1}-a-b\}$, with ${>,<,=,,}\odot\in\{$$\leq$$\geq\}$ denoting a comparator. Thereby, let
$L_{11}(r):=\textstyle\sum_{(a,b)\in\mathcal{S}_{a,b}}\sum_{(\lambda,\gamma)\in(S_{\lambda}\times\mathcal{S}_{\gamma})\cap\mathcal{S}_{<}}V,$
$L_{12}(r):=\textstyle\sum_{(a,b)\in\mathcal{S}_{a,b}}\sum_{(\lambda,\gamma)\in(S_{\lambda}\times\mathcal{S}_{\gamma})\cap\mathcal{S}_{>}}V,$
$L_{13}(r):=\textstyle\sum_{(a,b)\in\mathcal{S}_{a,b}}\sum_{(\lambda,\gamma)\in(S_{\lambda}\times\mathcal{S}_{\gamma})\cap\mathcal{S}_{=}}V.$
Then, the number of elementary events that lead to case ➁ is
$L_{2}=\Theta-(L_{11}+L_{12}+L_{13})$ (3)
where $L_{i}=\sum_{r=u+e_{1}-t}^{u+e_{1}+t}L_{i}(r)$ for $i\in\{a,b,c\}$.
In summary, for the cases in $\mathcal{L}$, we have
$\begin{cases}P_{{\textnormal{{EaED}}}}^{\textnormal{mc}}(R=r|u,e,e_{1})=(L_{11}(r)+L_{13}(r)/2)/\Theta,\\ P_{{\textnormal{{EaED}}}}^{\textnormal{fail}}(R=u+\frac{e}{2}|u,e,e_{1})=0,\\ P_{{\textnormal{{EaED}}}}^{\textnormal{succ}}(R=0|u,e,e_{1})=(L_{2}+L_{12}+L_{13}/2)/\Theta.\end{cases}$ (4)
Finally, we consider the cases in $\mathcal{M}$. First, note that they can only cause a miscorrection or a failure since $\bm{y}^{(1)},\bm{y}^{(2)}\notin\mathcal{S}_{t}({\bm{0}})$. In the following, we make a significant simplification by assuming that the outcomes of $\textsf{BDD}\left(\bm{y}^{(1)}\right)$ and $\textsf{BDD}\left(\bm{y}^{(2)}\right)$, whether miscorrection or failure, are independent of each other. In principle, this assumption is false as $d(\bm{y}^{(1)},\bm{y}^{(2)})=e$. However, as the positions of the erasures are random, we observe that fixing the decoding result of one BDD does not significantly influence the other one. Then, we obtain
$P_{{\textnormal{{EaED}}}}^{\textnormal{fail}}(u,e,e_{1})=P_{\textnormal{{BDD}}}^{\textnormal{fail}}(u\!+\!e_{1})P_{\textnormal{{BDD}}}^{\textnormal{fail}}(u\!+\!e\!-\!e_{1}).$ (5)
As for miscorrections, we assume that when both $\bm{y}^{(1)}$ and $\bm{y}^{(2)}$ are miscorrected, the probability of choosing one of the miscorrected codewords is $\frac{1}{2}$. Then we can approximate
$P_{{\textnormal{{EaED}}}}^{\textnormal{mc}}(r$ $|u,e,e_{1})=P_{\textnormal{{BDD}}}^{\textnormal{mc}}(r|u\!+\!e_{1})P_{\textnormal{{BDD}}}^{\textnormal{fail}}(u\!+\!e\!-\!e_{1})$
$\quad+P_{\textnormal{{BDD}}}^{\textnormal{fail}}(u\!+\!e_{1})P_{\textnormal{{BDD}}}^{\textnormal{mc}}(r|u\!+\!e\!-\!e_{1})$
$\quad+\tfrac{1}{2}P_{\textnormal{{BDD}}}^{\textnormal{mc}}(r|u\!+\!e_{1})P_{\textnormal{{BDD}}}^{\textnormal{mc}}(u\!+\!e\!-\!e_{1})$
$\quad+\tfrac{1}{2}P_{\textnormal{{BDD}}}^{\textnormal{mc}}(r|u\!+\!e\!-\!e_{1})P_{\textnormal{{BDD}}}^{\textnormal{mc}}(u\!+\!e_{1}).$ (6)
Now we have obtained the probabilities for cases in $\mathcal{L}$ and $\mathcal{M}$. Plugging the results into (A) and marginalizing over the decoding outcomes $\mathcal{D}$ yields Theorem 2.
Appendix B EaED with Anchor Bits
Based on the calculation in Appendix A, the DTP of the EaEDa can be obtained by considering that codewords found by $\textsf{BDD}\left(\bm{y}^{(1)}\right)$ and $\textsf{BDD}\left(\bm{y}^{(2)}\right)$ could be rejected by the MD. To this end, let $\bm{z}$ denote one of $\bm{y}^{(1)}$ or $\bm{y}^{(2)}$ and $\textsf{BDD}\left(\bm{z}\right)=\bm{c}$}. For $\bm{c}\in\mathcal{C}$}, define $n_{\textnormal{ca}}:=\{i\in[n]:z_{i}=0,c_{i}=1\}$} and $n_{\textnormal{wa}}:=\{i\in[n]:z_{i}=1,c_{i}=0\}$}. Then, the probability that $\bm{c}$ is accepted by the MD is $(1-P_{\textnormal{ca}})^{n_{\textnormal{ca}}}(1-P_{\textnormal{wa}})^{n_{\textnormal{wa}}}=:p_{\text{AC}}$, since the events of individual bits being anchor bits are statistically independent. Before we derive the $P_{{\textnormal{{EaED}}}^{\textsf{a}}}$, we first compute the DTP of decoding one random test pattern, say, $\bm{y}^{(1)}$ when $u+e_{1}>t$. It is denoted as $P_{\textnormal{{EaEDs}}^{\textsf{a}}}$. We define
$L_{s}(r)=\sum_{a,b,n_{\textnormal{ca}},n_{\textnormal{wa}}}A_{r}\tbinom{r}{a}\tbinom{n-r}{b}\tbinom{b}{n_{\textnormal{wa}}}\tbinom{a}{n_{\textnormal{ca}}}\tbinom{r-a}{c}\tbinom{n-r-b}{d}\cdot p_{\text{AC}}$
where $c=e_{1}-(b-n_{\textnormal{wa}})$ and $d=e-e_{1}-(a-n_{\textnormal{ca}})$. Then, the miscorrection probability is
$P_{\textnormal{{EaEDs}}^{\textsf{a}}}^{\textnormal{mc}}(r|u,e,e_{1})=L_{s}(r)/\Theta,$ (7)
and $P_{\textnormal{{EaEDs}}^{\textsf{a}}}^{\textnormal{fail}}(u,e,e_{1})=1-P_{\textnormal{{EaEDs}}^{\textsf{a}}}^{\textnormal{mc}}(u,e,e_{1})$ as the success probability is 0 when $u+e_{1}>t$.
Next, we analyze the decoding result by cases.
Case 1: For $e=0$, EaEDa reverts to BDDa.
Case 2: For $e\neq 0$, we need to compute two cases.
Case 2.1: If $2u+e<d_{\textnormal{min}}$, EaED is guaranteed to succeed. However, due to the MD step, there are three sub-cases.
Case 2.1.1: If $u\!+\!e_{1}\leq t$} and $u\!+\!e\!-\!e_{1}<d_{\textnormal{min}}\!-\!t$} or if $u\!+\!e_{1}<d_{\textnormal{min}}\!-\!t$} and $u\!+\!e\!-\!e_{1}\leq t$}, either $\bm{y}^{(1)}$ or $\bm{y}^{(2)}$ will be decoded to $\bm{c}=\bm{0}$ (with $n_{\textnormal{wa}}=u$ and $n_{\textnormal{ca}}=0$) and the other one results in a decoding failure. Let $P_{\bm{0}}:=(1-P_{\textnormal{wa}})^{u}$. Then $P_{{\textnormal{{EaED}}^{\textsf{a}}}}^{\textnormal{succ}}(u,e)=P_{\bm{0}}$} and $P_{{\textnormal{{EaED}}^{\textsf{a}}}}^{\textnormal{fail}}(u,e)=1-P_{\bm{0}}$}.
Case 2.1.2: If $u+e_{1}\leq t$ and $u+e-e_{1}\geq d_{\min}-t$. $\textsf{BDD}\left(\bm{y}^{(1)}\right)\!=\!\bm{0}$}, and $\textsf{BDD}\left(\bm{y}^{(2)}\right)$ either fails or miscorrects. For a miscorrection where $\textsf{BDD}\left(\bm{y}^{(2)}\right)=\bm{c}\neq\bm{0}$ with $\textnormal{wt}\left(\bm{c}\right)=r$, $d_{\sim{\textnormal{E}(\bm{y})}}(\bm{0})<d_{\sim{\textnormal{E}(\bm{y})}}(\bm{c})$ always holds, as shown in [7, Appendix A]. Therefore, if $\bm{0}$ is accepted by the MD (probability $P_{\bm{0}}$), it will be chosen as the EaED decision. Then, $P_{{\textnormal{{EaED}}^{\textsf{a}}}}^{\textnormal{succ}}(u,e)=P_{\bm{0}}$}. When $\bm{0}$ is rejected (probability $(1-P_{\bm{0}})$), EaED will either fail or miscorrect. We compute $P_{\textnormal{{EaED}}^{\textsf{a}}}^{\textnormal{mc}}(r|u,e)\!=\!(1-P_{\bm{0}})P_{\textnormal{{EaEDs}}^{\textsf{a}}}^{\textnormal{mc}}\left(r|u,e,e_{2}\right)$ and $P_{{\textnormal{{EaED}}^{\textsf{a}}}}^{\textnormal{fail}}(u,e)=1-P_{{\textnormal{{EaED}}^{\textsf{a}}}}^{\textnormal{succ}}(u,e)-P_{\textnormal{{EaED}}^{\textsf{a}}}^{\textnormal{mc}}(u,e)$.
Case 2.1.3: If $u+e-e_{1}\leq t$ and $u+e_{1}\geq d_{\min}-t$, the probabilities are calculated with the same formulas as in case 2.1.2, by redefining $d_{\bm{y}\bm{c}}=d_{\min}-(u+e_{1})$.
The case where both $\bm{y}^{(1)}$ and $\bm{y}^{(2)}$ are miscorrected is impossible as it contradicts the condition $2u+e<d_{\textnormal{min}}.$ Next, we consider the case where $2u+e\geq d_{\textnormal{min}}$.
Case 2.2: $2u+e\geq d_{\textnormal{min}}$. Analogous to Appendix A, we divide the cases into sets $\mathcal{L}$, $\mathcal{M}$, and $\mathcal{R}$, and observe that we can replace the probabilities associated with EaED in (A) by their EaEDa counterparts.
For set $\mathcal{L}$, as depicted in Fig. 5, we write $P_{\bm{0}}=(1-P_{\textnormal{wa}})^{u}$ and $P_{\bm{c}}=(1-P_{\textnormal{ca}})^{a-(e_{1}-\lambda)}(1-P_{\textnormal{wa}})^{b-\gamma}$. We can reuse (4), replacing $L_{11}$, $L_{12}$, and $L_{13}$ by $L_{11}^{\mathsf{a}}$, $L_{12}^{\mathsf{a}}$, and $L_{13}^{\mathsf{a}}$, respectively, which are calculated as
$L_{11}^{\mathsf{a}}=\sum_{a,b}\bigg(\sum_{(\lambda,\gamma)\in(S_{\lambda}\times\mathcal{S}_{\gamma})\cap\mathcal{S}_{<}}\!\!\!\!\!\!\!\!\!\!\!\!V\cdot P_{\bm{c}}+\sum_{(\lambda,\gamma)\in(S_{\lambda}\times\mathcal{S}_{\gamma})\cap\mathcal{S}_{\geq}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!V\cdot(1-P_{\bm{0}})P_{\bm{c}}\bigg)$
$L_{12}^{\mathsf{a}}=\sum_{a,b}\bigg(\sum_{(\lambda,\gamma)\in(S_{\lambda}\times\mathcal{S}_{\gamma})\cap\mathcal{S}_{\leq}}\!\!\!\!\!\!\!\!\!\!\!\!\!V\cdot P_{\bm{0}}(1-P_{\bm{c}})+\sum_{(\lambda,\gamma)\in(S_{\lambda}\times\mathcal{S}_{\gamma})\cap\mathcal{S}_{>}}\!\!\!\!\!\!\!\!\!\!\!\!\!V\cdot P_{\bm{0}}\bigg)$
$L_{13}^{\mathsf{a}}=\sum_{a,b}\sum_{(\lambda,\gamma)\in(S_{\lambda}\times\mathcal{S}_{\gamma})\cap\mathcal{S}_{=}}\!\!\!\!\!\!\!\!\!\!\!\!\!\!V\cdot P_{\bm{0}}P_{\bm{c}}$
where $(a,b)\in\mathcal{S}_{a,b}$. Additionally, define
$L_{\mathsf{fail}}=\textstyle\sum_{(a,b)\in\mathcal{S}_{a,b}}\sum_{(\lambda,\gamma)\in(S_{\lambda}\times\mathcal{S}_{\gamma})}V(1-P_{\bm{c}})(1-P_{\bm{0}}).$
Then, we have
$\begin{cases}P_{{\textnormal{{EaED}}^{\textsf{a}}}}^{\textnormal{mc}}(R=r|u,e,e_{1})\!=\!(L_{11}^{\mathsf{a}}(r)+L^{\mathsf{a}}_{13}(r)/2)/\Theta,\\ P_{{\textnormal{{EaED}}^{\textsf{a}}}}^{\textnormal{fail}}(u,e,e_{1})=(L_{\mathsf{fail}}+L_{2}(1-P_{\bm{0}}))/\Theta,\\ P_{{\textnormal{{EaED}}^{\textsf{a}}}}^{\textnormal{succ}}(u,e,e_{1})\!=\!(L_{2}P_{\bm{0}}+L^{\mathsf{a}}_{12}+L^{\mathsf{a}}_{13}/2)/\Theta.\end{cases}$ (8)
Finally, we consider the set $\mathcal{M}$, where we still assume that the decoding results of $\bm{y}^{(1)}$ and $\bm{y}^{(2)}$ are independent. Then, the probabilities are obtained in the same way as for EaED in (A), except that the BDD probabilities are replaced with those of a single test-pattern EaED computed by (7). Finally, the DTP of EaEDa can be computed using the framework of Theorem 2.

### v2

@@
 However, BDD falls short in many applications.
-In this work, we consider error-and-erasure decoding and its variants that improve upon BDD.
+In this work, we consider error-and-erasure decoding and its improved variants.
 We derive closed-form expressions for their error probabilities and validate them through simulations.
@@
 With the DTP, the post-decoding bit error rate (BER) and frame error rate (FER) for various memoryless channels can be computed.
-The DTP is validated through simulations and applied in a concatenated coding scheme.
+The DTP is validated through simulations and applied in a concatenated scheme.
 […]
@@
 The exact weight enumerators of some short BCH codes are listed in [9].
+Note that both [8] and the results derived in this work rely on the weight distribution.
+If only an approximate weight distribution is available, the resulting analysis is also approximate.
 We consider three communication channels: the binary symmetric channel (BSC) with cross-over probability $\delta=P(Y\neq X)$}, the EaE channel, and the binary input additive white Gaussian noise (BI-AWGN) channel.
@@
 Thus, the parameters of the resulting EaE are $\delta_{c}=Q\left(\tfrac{\mathsf{T}+1}{\sigma}\right)$, and $\epsilon_{c}=1-Q\left(\tfrac{\mathsf{T}-1}{\sigma}\right)-\delta_{c}.$ For $\mathsf{T}=0$, we recover a BSC with $\delta=Q\left(\tfrac{1}{\sigma}\right)$.
-We use $\bm{x}\in\mathcal{C}$ to denote the transmitted codeword and $\bm{y}$ to denote the received word, whose alphabet is determined by the channel.
-At the output of a BSC, let $\mathcal{S}_{t}(\bm{c})=\{\bm{y}\in\mathbb{F}_{2}^{n}:d(\bm{y},\bm{c})\leq t\}$ denote the Hamming sphere of radius $t$ for $\bm{c}\in\mathcal{C}$.
-Then
-$\textsf{BDD}\left(\bm{y}\right)=\begin{cases}\bm{c},&\text{if }\bm{y}\in\mathcal{S}_{t}(\bm{c}),\\ \bm{y}\,(\mathsf{fail}),&\text{otherwise}.\end{cases}$
-Note that the decoding outcome falls into one of three categories: (i) decoding success, if $\bm{c}=\bm{x}$; (ii) decoding failure if no codeword is found; or (iii) miscorrection (also called a decoding error as in [8]), if $\bm{c}\neq\bm{x}$.
+We use $\bm{x}\in\mathcal{C}$ to denote the transmitted codeword and $\bm{y}$ to denote the received word, whose alphabet is determined by the channel. At the output of a BSC, we have
+$\textsf{BDD}\left(\bm{y}\right)=\begin{cases}\bm{c},&\text{if }\bm{y}\in\mathcal{S}_{t}(\bm{c}),\\ \bm{y}\,(\mathsf{fail}),&\text{otherwise},\end{cases}$
+where $\mathcal{S}_{t}(\bm{c})=\{\bm{y}\in\mathbb{F}_{2}^{n}:d(\bm{y},\bm{c})\leq t\}$ denote the Hamming sphere of radius $t$ for $\bm{c}\in\mathcal{C}$.
+The decoding outcome falls into one of three categories: (i) decoding success, if $\bm{c}=\bm{x}$; (ii) decoding failure if no codeword is found; or (iii) miscorrection (also called a decoding error as in [8]), if $\bm{c}\neq\bm{x}$.
 For each decoder considered in this paper, we distinguish these three cases when computing the DTP.
@@
 $P_{\textnormal{{BDD}}}(r|u)=P_{\textnormal{{BDD}}}^{\textnormal{mc}}(r|u)+P_{\textnormal{{BDD}}}^{\textnormal{fail}}(r|u)+P_{\textnormal{{BDD}}}^{\textnormal{succ}}(r|u).$
-Note that for a decoding failure, $P_{\textnormal{{BDD}}}^{\textnormal{fail}}(R\neq u|u)=0$, since $\textsf{BDD}\left(\bm{y}\right)=\bm{y}$ in this case, and for a decoding success, $P_{\textnormal{{BDD}}}^{\textnormal{succ}}(R\neq 0|u)=0$.
-Furthermore, we denote by $P_{\textnormal{{BDD}}}^{\textnormal{succ}}(u)$, $P_{\textnormal{{BDD}}}^{\textnormal{fail}}(u)$, and $P_{\textnormal{{BDD}}}^{\textnormal{mc}}(u)$ the conditional probabilities of success, failure, and miscorrection given $u$ input errors, respectively.
+Note that for a decoding failure where $\textsf{BDD}\left(\bm{y}\right)=\bm{y}$}, $P_{\textnormal{{BDD}}}^{\textnormal{fail}}(R\neq u|u)\!=\!0$}, while for a decoding success, $P_{\textnormal{{BDD}}}^{\textnormal{succ}}(R\neq 0|u)\!=\!0$}.
+We further denote by $P_{\textnormal{{BDD}}}^{\textnormal{succ}}(u)$, $P_{\textnormal{{BDD}}}^{\textnormal{fail}}(u)$, and $P_{\textnormal{{BDD}}}^{\textnormal{mc}}(u)$ the conditional probabilities of success, failure, and miscorrection, respectively, given $u$ errors.
 Since we study linear codes with a symmetric decoding algorithm, we assume the transmission of $\bm{x}=\bm{0}$} throughout the paper.
@@
 Additionally, the number of codewords $\bm{c}\in\mathcal{C}$ with $\textnormal{wt}\left(\bm{c}\right)=r,$ is given by the weight enumerator $A_{r}$.
-Thus, we obtain
+Thus,
 $P_{\textnormal{{BDD}}}^{\textnormal{mc}}(r|u)=\frac{N(u)}{\tbinom{n}{u}}=\frac{\sum_{(a,b)\in\mathcal{S}_{a,b}}A_{r}\tbinom{r}{a}\tbinom{n-r}{b}}{\tbinom{n}{u}}$.
@@
 For $2u+e\geq d_{\textnormal{min}}$}, $P_{\textnormal{{EaED}}}(r|u,e)=\sum_{d\in\mathcal{D}}P_{\textnormal{{EaED}}}^{d}(r|u,e)$ and
-$P_{\textnormal{{EaED}}}^{d}(r|u,e)=\textstyle\sum_{e_{1}=0}^{e}P_{\textnormal{{EaED}}}^{d}(r|u,e,e_{1})\cdot P(e_{1}|u,e),$
+$P_{\textnormal{{EaED}}}^{d}(r|u,e)=\textstyle\sum_{e_{1}=0}^{e}P_{\textnormal{{EaED}}}^{d}(r|u,e,e_{1})\cdot P(e_{1}|u,e),$ (1)
 where $e_{1}=\!\textnormal{wt}\left(\bm{p}^{(1)}\right)$}.
@@
 The OSD uses order 2, as higher orders do not yield additional performance gains.
-Using the $[255,239,5]$ BCH code as an example, Tab.
-I and Tab.
-II list the conditional probabilities $P_{{\textnormal{{EaED}}}}^{\textnormal{succ}}(u,e)$ and $P_{{\textnormal{{EaED}}}}^{\textnormal{mc}}(u,e)$, respectively, which are calculated using (2) and marginalization over $U$.
+Using the $[255,239,5]$ BCH code as an example, Tables I and II list the conditional probabilities $P_{{\textnormal{{EaED}}}}^{\textnormal{succ}}(u,e)$ and $P_{{\textnormal{{EaED}}}}^{\textnormal{mc}}(u,e)$, respectively.
+These quantities represent the probabilities of successful decoding and miscorrection given $u$ errors and $e$ erasures, computed using (1) and marginalized over $R$.
 Note that if $e=0$, Theorem 2 reverts to Theorem 1.
@@
 II, reflecting that EaED has a stronger error correction capability but is more prone to miscorrection.
-By contrast, EaEDa substantially reduces miscorrections through the use of anchor bits.
+By contrast, EaEDa substantially reduces miscorrections through the use of anchor bits, which can be seen from the dashed lines in Fig.
+2.
 Thus, EaEDa improves decoding performance while maintaining a low miscorrection rate, keeping the undetectable error rate negligible.
@@
 However, as the positions of the erasures are random, we observe that fixing the decoding result of one BDD does not significantly influence the other one.
-Then, we obtain
+Thus,
 $P_{{\textnormal{{EaED}}}}^{\textnormal{fail}}(u,e,e_{1})=P_{\textnormal{{BDD}}}^{\textnormal{fail}}(u\!+\!e_{1})P_{\textnormal{{BDD}}}^{\textnormal{fail}}(u\!+\!e\!-\!e_{1}).$ (6)
@@
 Finally, the DTP of EaEDa can be computed using the framework of Theorem 2.
+Acknowledgment
+This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101001899).